Please help find the Standard Error and the last question about confidence interval

Durable press cotton fabrics are treated to improve their recovery from wrinkles after washing. Wrinkle recovery angle measures how well a fabric recovers from wrinkles. Higher is better. Here are data on the wrinkle recovery angle (in degrees) tor the same fabric specimens discussed ?n the previous exercise Permafresh Hylite 131 142 108 130 137 160 149 158 190 A manufacturer wants to know how large is the difference in mean wrinkle recovery angle a. The mean and the standard deviation for the Permafresh are ? 1129.600 Round to 3 decimal places 13.012 Round to 3 decimal places b. The mean and the standard deviation for the Hylite are 164.250 Round to 3 decimal places s-17.821 Round to 3 decimal places c. The standard error is SE -34.65 Round to 3 decimal places

"Durable press" cotton fabrics are treated to improve their recovery from wrinkles after washing. "Wrinkle recovery angle" measures how well a fabric recovers from wrinkles. Higher is better. Here are data on the wrinkle recovery angle (in degrees) for the same fabric specimens discussed in the previous exercise:

A manufacturer wants to know how large is the difference in mean wrinkle recovery angle.

a. The mean and the standard deviation for the Permafresh are:

b. The mean and the standard deviation for the Hylite are:

c. The standard error is:

SE=-34.65 Round to 3 decimal places.

The critical value from the distribution for a confidence interval of 99.8% is:

d. Using the the answer from above (all rounded to 3 decimal places), calculate by hand a 99.8% confidence interval for the difference in mean wrinkle recovery angle:
Answer #1

From the given information, for Permafresh: \(\bar{x}_{1}=129.60, s_{1}=13.01, n_{1}=5\), for Hylite: \(\bar{x}_{2}=8.91, s_{2}=17.82\), and \(n_{2}=4\)

Standard Error is \(S_{\left(\pi_{1}-x_{2}\right)}=\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}=\sqrt{\frac{13.01^{2}}{5}+\frac{17.82^{2}}{4}}=10.6414\)

\(100(1-\alpha) \%\) confidence interval for difference of true means is \(\left(\left(\bar{x}_{1}-\bar{x}_{2}\right)-t_{\alpha / 2} \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}},\left(\bar{x}_{1}-\bar{x}_{2}\right)+t_{\alpha / 2} \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}\right)\)

\(99.8 \%\) confidence interval for the difference in mean wrinkle recovery angle is \(\left(\left(\bar{x}_{1}-\bar{x}_{2}\right)-t_{0.01 / 2} \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}},\left(\bar{x}_{1}-\bar{x}_{2}\right)+t_{0.01 / 2} \sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}\right)\)

\(=\left((129.6-164.3)-3.499 \sqrt{\frac{(13)^{2}}{5}+\frac{(17.8)^{2}}{4}},(129.6-164.3)+3.499 \sqrt{\frac{(13)^{2}}{5}+\frac{(17.8)^{2}}{4}}\right)\)

\(=(-71.8965,2.4965)\)

\(99.8 \%\) confi dence interval is \(-71.8965 \leq \mu_{1}-\mu_{2} \leq 2.4965 .\)